### v1

@@
Energy transport in holographic
non-conformal interfaces
@@ 1 Introduction
It has been established in [3] the transmission coefficient $\mathcal{T}$ and reflection coefficient $\mathcal{R}$ fall within the interval $[0,1]$ for conformal interfaces in a class of unitary CFTs, while they break these bounds for nonunitary CFTs. The relation $\mathcal{T}+\mathcal{R}=1$ always holds. For ICFT, similar results were derived from averaged null-energy condition [4]. In strongly coupled systems, the AdS/CFT correspondence proves useful [10, 11]. For holographic ICFTs, which are dual to a gravitational system consists of two AdS3 regions with the same cosmological constant separated by an interface brane with tension $\sigma$, the transmission coefficient is given by $\mathcal{T}=\left(1+4\pi GL\sigma\right)^{-1}$ and $\mathcal{R}=1-\mathcal{T}$ [12]. These findings were further confirmed in a study of holographic non-equilibrium steady states with an interface brane [13]. Given that the tension $\sigma$ is constrained to $[0,1/(4\pi GL)]$ in AdS/ICFT, new constraints $\frac{1}{2}\leq\mathcal{T}\leq 1\,,$ were discovered [12]. The permissible range of transmission coefficient can be further tunned from 00 to $1$ in double brane holographic models [14] and Janus configurations [15]. Further studies on the energy transports in AdS/ICFT can be found in [16, 17, 18, 19, 20].
@@ 2.1 Solution at first order and energy transport
~$\begin{split}\tilde{\mathcal{R}}&=\frac{\text{reflected energy flux}}{\text{injected energy flux}}=\frac{T_{\text{re}}}{T_{\text{in}}}=\frac{|\mathcal{R}|\cos(\omega t+\phi_{r})}{\cos(\omega t)}\,,\\ \tilde{\mathcal{T}}&=\frac{\text{transmitted energy flux}}{\text{injected energy flux}}=\frac{T_{\text{tr}}}{T_{\text{in}}}=\frac{|\mathcal{T}|\cos(\omega t+\phi_{t})}{\cos(\omega t)}\,.\end{split}$ (2.12)
@@
~$\tilde{\mathcal{R}}+\tilde{\mathcal{T}}=1\,.$ (2.13)
@@
    ~$\tilde{\mathcal{R}}=\mathcal{R}\,,~{}~{}~{}\tilde{\mathcal{T}}=\mathcal{T}\,.$ (2.14)
@@
  • The instantaneous values of ~$\tilde{\mathcal{R}},\tilde{\mathcal{T}}$ depend on the phase $\omega t$ of the injected wave at the interface. When $\omega t=(n+\frac{1}{2})\pi$ for integer $n$, at the interface the injected wave $\langle T_{\text{in}}\rangle=2\epsilon\cos(\omega t)=0$, and $\langle T_{\text{re}}\rangle=\langle T_{\text{tr}}\rangle$.
@@
    where ,′′$\mathcal{R}^{\prime},\mathcal{R}^{\prime\prime},\mathcal{T}^{\prime},\mathcal{T}^{\prime\prime}$ are real and satisfy
@@
    ~=′′$\tilde{\mathcal{R}}=\mathcal{R}^{\prime}-\mathcal{R}^{\prime\prime}\tan(\omega t)\,,~{}~{}~{}\tilde{\mathcal{T}}=\mathcal{T}^{\prime}-\mathcal{T}^{\prime\prime}\tan(\omega t)\,.~{}~{}~{}$ (2.17)
    We define the averaged transmission and reflection coefficients by performing the principle integration from $-\frac{\pi}{2\omega}$ to $\frac{\pi}{2\omega}$,
    ~$\begin{split}\langle\tilde{\mathcal{R}}\rangle=\mathcal{R}^{\prime}\,,~{}~{}~{}~{}\langle\tilde{\mathcal{T}}\rangle=\mathcal{T}^{\prime}\,.~{}~{}~{}~{}\end{split}$ (2.18)
@@ 2.1.1 The equations of perturbations
$\begin{split}&t_{\text{I}}=t+\frac{G\epsilon}{L}e^{i\omega\tau}\lambda(z)+c.c.\,,\ x_{\text{I}}=p(z)+\frac{2G\epsilon}{L}e^{i\omega\tau}h_{\text{I}}(z)+c.c.\,,\ u_{\text{I}}=z+\frac{2G\epsilon}{L}e^{i\omega\tau}\xi(z)+c.c.\,,\\ &t_{\text{II}}=t-\frac{G\epsilon}{L}e^{i\omega\tau}\lambda(z)+c.c.\,,\ x_{\text{II}}=-p(z)+\frac{2G\epsilon}{L}e^{i\omega\tau}h_{\text{II}}(z)+c.c.\,,\ u_{\text{II}}=z-\frac{2G\epsilon}{L}e^{i\omega\tau}\xi(z)+c.c.\,,\end{split}$ (2.20)
@@ 3.1 On the UV boundary condition for asymptotic AdS2 solution
where $A\not=0$ and $B\leq 0$. For a given profile of $Q$, the scalar field $\phi(z)$ and the potential $V(\phi)$ are determined from Eq. (2.5). The localized scalar operator in the dual ICFT has a conformal dimension of $1/2$ for $B<0$ and $1$ for $B=0$. Below, we show that under the sourceless boundary condition for the fluctuations, the relation $\mathcal{R}+\mathcal{T}=1$ holds, consistent with energy conservation as discussed in Sec. 2.1.
[…]
$\xi(z)=\xi_{1}\,z+\xi_{2}\,z^{2}+\frac{1}{2}\left[-1-\mathcal{R}+\mathcal{T}-\omega^{2}(1+A^{2})\xi_{1}+\frac{2B\xi_{2}}{A}\right]z^{3}+\cdots\,.$ (3.2)
@@
where $p_{1}$ and $p_{2}$ are two integral constants. For $B=0$, the UV expansion is
@@
where $p_{1}$ and $p_{2}$ are again integration constants. For $B<0$, applying the sourceless condition $p_{1}=0$ yields, $p(z)=p_{2}\sqrt{z}\left(1+\mathcal{O}(z)\right).$ For $B=0$, the sourceless condition gives $p(z)=p_{2}\,z+\cdots$. Note that in (3.1) if the nonzero higher order term of $z^{n}$ first starts with an integer $n$ of $n\geq 2$, the scalar fluctuation have higher conformal dimension $n-1$.
@@
when $z\to 0$ By imposing the infalling boundary condition for $\xi_{l}$ near the IR and the above UV boundary condition, we can solve the system and obtain the final solution for $\xi_{l}$ and the transmission coefficient $\mathcal{T}$.
@@ 3.2 An example from numerical studies
where the interface entropy is defined in [21], and $c=3L/2G$ is the central charge of the left/right CFT. We restrict our discussion to the region $b>0$ to ensure a positive interface entropy. The UV expansion of this profile matches exactly with Eq. (3.1) upon noticing that $A=a$ and $B=b-a$.
@@
In Fig. 2, we show the plots of $\mathcal{T}^{\prime}$ and $\phi_{t}$ as functions of $a$. In the upper two plots, we fix the frequency and tune $b$. Both $\mathcal{T}^{\prime}$ and $\phi_{t}$ exhibit oscillations with a period of approximately $\omega a/\langle\mathcal{O}\rangle^{2}\simeq 0.4$. The amplitude of these oscillation decreases as the parameter $b$ is reduced or $\omega/\langle\mathcal{O}\rangle^{2}$ is increased.
Figure 3: The plots of $\mathcal{T}^{\prime}$ and $\phi_{t}$ as functions of $b$, for different values of $a$ with fixed $\omega/\langle\mathcal{O}\rangle^{2}=0.01$ (upper two plots) and for different values of $\omega/\langle\mathcal{O}\rangle^{2}$ with fixed $a=2$ (lower two plots). In the upper two plots, $a=1$ (black), $3$ (blue), $5$ (purple), and $7$ (red).
In Fig. 3, we show the plots of $\mathcal{T}^{\prime}$ and $\phi_{t}$ as functions of $b$. Due to the null energy condition constraint $a\geq b$, we focus on the region where $b\in(0,1]$ with $a\geq 1$. We observe that, for the parameters considered, $\mathcal{T}^{\prime}$ and $\phi_{t}$ exhibit monotonic behavior for all values of $b$ within this region. This behavior contrasts with the trends observed in Fig. 2. Moreover, we find that when $b\to 0$, $\mathcal{T}^{\prime}$ approaches distinct values, which differs from its dependence on $a$. Furthermore, for a fixed value of $b$, we observe that $\mathcal{T}^{\prime}$ increases when $\omega/\langle\mathcal{O}\rangle^{2}$ decreases, and this is consistent with the observations in Fig. 1.
3.3 Two analytical examples
In this subsection we show two brane profiles for which analytical solution for the transmission coefficients can be derived. The induced metric for these two examples are asymptotically AdS2 in the both UV and IR region. From the analytical solution, we observe the key features that were previously identified in the numerical example in Sec. 3.2.
3.3.1 The solution I with UV and IR asymptotically AdS2
@@
The null energy condition requires that $ab>0$. Here we consider $a>0$ and $b>0$. The background brane geometry is asymptotically AdS with radius $L\sqrt{1+a^{2}}$ in the UV and radius $L$ in the IR. From the UV expansion, it exactly matches with (3.1) by identifying $A=a$ and $B=-ab/2$.
@@
$\begin{split}\xi_{l}(z)=&\,~{}c_{-}\,ze^{-i\sqrt{2}\omega z}(bz+1)^{\frac{ia\omega}{b}}\,U\left(\frac{1}{2}-\frac{i(\sqrt{2}-4a)\omega}{4b},~{}1+\frac{2ia\omega}{b},~{}\frac{2i\sqrt{2}(1+bz)\omega}{b}\right)\,\\ &+c_{+}\,ze^{-i\sqrt{2}\omega z}(bz+1)^{\frac{ia\omega}{b}}L\,\left(-\frac{1}{2}+\frac{i(\sqrt{2}-4a)\omega}{4b},~{}\frac{2ia\omega}{b},~{}\frac{2i\sqrt{2}(1+bz)\omega}{b}\right)\,,\end{split}$ (3.20)
@@
[…]
@@
3.3.2 The solution II with UV and IR asymptotically AdS2
We consider another brane profile
$\psi(z)=\frac{2a}{b}(\sqrt{1+bz}-1)\,.$ (3.25)
The null energy condition requires $ab>0$ and we again assume $a>0$ and $b>0$. The background brane geometry is asymptotically AdS with radius $\sqrt{1+a^{2}}L$ in the UV and $L$ in the IR. From the UV expansion, it exactly matches (3.1) by identifying that $A=a,B=-ab/4$.
The profile for the scalar field is
[…]
and the potential is
$V(\phi(z))=\frac{a(4+4a^{2}+5bz)}{2L(1+a^{2}+bz)^{3/2}}\,.$ (3.27)
The solution to the homogeneous part of ODE (2.22) takes the form
$\begin{split}\xi_{l}(z)=&\,~{}c_{-}e^{-i\sqrt{\frac{3}{2}}\omega z}z\sqrt{1+bz}U\left(\frac{3}{4}+\frac{i(2a^{2}-1)\omega}{2\sqrt{6}b},\,\frac{3}{2},\,\frac{i\sqrt{6}(1+bz)\omega}{b}\right)\\ &+c_{+}e^{-i\sqrt{\frac{3}{2}}\omega z}z\sqrt{1+bz}L\left(-\frac{3}{4}-\frac{i(2a^{2}-1)\omega}{2\sqrt{6}b},\,\frac{1}{2},\,\frac{i\sqrt{6}(1+bz)\omega}{b}\right)\,,\end{split}$ (3.28)
From the solution (3.28) with $c_{+}=0$, we obtain the transmission coefficient
$\mathcal{T}=1+\frac{4iab\omega}{C_{2}}U\left(\frac{3}{4}+\frac{i(2a^{2}-1)\omega}{2\sqrt{6}b},~{}\frac{3}{2},~{}\frac{i\sqrt{6}\omega}{b}\right)\,,$ (3.29)
where
$\begin{split}C_{2}=&\,~{}2b\left[b-i(2a+\sqrt{6})\omega\right]U\left(\frac{3}{4}+\frac{i(2a^{2}-1)\omega}{2\sqrt{6}b},~{}\frac{3}{2},~{}\frac{i\sqrt{6}\omega}{b}\right)\\ &+(4a^{2}\omega-2\omega-3i\sqrt{6}b)\omega\,U\left(\frac{7}{4}+\frac{i(2a^{2}-1)\omega}{2\sqrt{6}b},~{}\frac{5}{2},~{}\frac{i\sqrt{6}\omega}{b}\right)\,,\end{split}$ (3.30)
and
$\mathcal{R}=1-\mathcal{T}\,,\ \ c_{-}=-\frac{8iab}{\omega C_{2}}\,.$ (3.31)
@@ 3.4 Another kind of UV asymptotically AdS geometry
For the specific case $\psi(z)=-az^{2}$, the exact solution for $\xi(z)$ exists. The infalling boundary condition and $\xi_{1}=0$ imply $\xi_{3}=0$, leading to $\mathcal{R}=0,\mathcal{T}=1$.
@@ A.2 Equations for the fluctuations
$\begin{split}\lambda=&\,~{}\frac{(\bm{I}+\bm{R}-\bm{T})z^{3}+2\xi}{i\omega z}\,,\\ h_{\text{I}}=&\,~{}\frac{(1-\omega^{2}z^{2})\xi-z\xi^{\prime}}{\omega^{2}\psi^{\prime}z^{2}}+\sqrt{\frac{-Lz}{\psi^{\prime\prime}}}\frac{(1+\psi^{\prime 2})\left(\psi^{\prime\prime}+z\psi^{(3)}\right)-z\psi^{\prime}\psi^{\prime\prime}}{\omega^{2}(1+\psi^{\prime 2})^{7/4}z}p\,\\ &+\frac{2\sqrt{-Lz\psi^{\prime\prime}}}{\omega^{2}(1+\psi^{\prime 2})^{3/4}}p^{\prime}+\frac{1}{\omega^{3}\psi^{\prime}(1+\psi^{\prime 2})}\cdot\text{\bf s}\,,\\ h_{\text{II}}=&\,-h_{\text{I}}+\frac{2(1-\omega^{2}z^{2})\xi-2z\xi^{\prime}}{\omega^{2}\psi^{\prime}z^{2}}-\frac{z\left[2(\bm{I}+\bm{R}-\bm{T})+i(\bm{I}-\bm{R}+\bm{T})\omega\psi^{\prime}z\right]}{\omega^{2}\psi^{\prime}}\,,\end{split}$ (A.7)

### v2

@@
We study energy transport in a system of two dimensional conformal field theories exchanging energy across a non-conformal interface involving a localised scalar operator, using holographic duality. By imposing the sourceless boundary condition, or equivalently, enforcing energy conservation at the interface, we show that the sum of the transmission and reflection coefficients is equal to one. Unlike conformal interfaces, we find that both the energy transmission and reflection coefficients are generally complex and frequency dependent. When the induced metric on the interface brane connects two distinct AdS2 geometries, the transmission coefficient approaches the value expected for a conformal interface in the UV regime at high frequencies and in the IR regime at low frequencies. In the intermediate frequency range, the transmission coefficient may exhibit oscillatory behavior. Moreover, we present a nontrivial example of a fully transmissive interface, which exhibits similarities to a topological interface.
@@ 1 Introduction
It has been established in [3] the transmission coefficient $\mathcal{T}$ and reflection coefficient $\mathcal{R}$ fall within the interval $[0,1]$ for conformal interfaces in a class of unitary CFTs, while they break these bounds for nonunitary CFTs. The relation $\mathcal{T}+\mathcal{R}=1$ always holds. For ICFT, similar results were derived from averaged null-energy condition [4]. In strongly coupled systems, the AdS/CFT correspondence proves useful [10, 11]. For holographic ICFTs, which are dual to a gravitational system consists of two AdS3 regions with the same cosmological constant separated by an interface brane with tension $\sigma$, the transmission coefficient is given by $\mathcal{T}=\left(1+4\pi GL\sigma\right)^{-1}$ and $\mathcal{R}=1-\mathcal{T}$ [12]. These findings were further confirmed in a study of holographic non-equilibrium steady states with an interface brane [13]. Given that the tension $\sigma$ is constrained to $0,1/(4\pi GL)]$ in AdS/ICFT, new constraints $\frac{1}{2}\leq\mathcal{T}\leq 1\,$, were discovered [12]. The permissible range of transmission coefficient can be further tunned from 0 to $1$ in double brane holographic models [14] and Janus configurations [15]. Further studies on the energy transports in AdS/ICFT can be found in [16, 17, 18, 19, 20].
@@ 2.1 Solution at first order and energy transport
[…]
@@
$\tilde{\mathcal{R}}+\tilde{\mathcal{T}}=1\,.$ (2.13)
@@
    $\tilde{\mathcal{R}}=\mathcal{R}\,,~~~\tilde{\mathcal{T}}=\mathcal{T}\,.$ (2.14)
@@
  • The instantaneous values of $\tilde{\mathcal{R}},\tilde{\mathcal{T}}$ depend on the phase $\omega t$ of the injected wave at the interface. When $\omega t=(n+\frac{1}{2})\pi$ for integer $n$, at the interface the injected wave $\langle T_{\text{in}}\rangle=2\epsilon\cos(\omega t)=0$, and $\langle T_{\text{re}}\rangle=\langle T_{\text{tr}}\rangle$.
@@
    where $\mathcal{R}^{\prime},\mathcal{R}^{\prime\prime},\mathcal{T}^{\prime},\mathcal{T}^{\prime\prime}$ are real and satisfy
@@
    $\tilde{\mathcal{R}}=\mathcal{R}^{\prime}-\mathcal{R}^{\prime\prime}\tan(\omega t)\,,~~~\tilde{\mathcal{T}}=\mathcal{T}^{\prime}-\mathcal{T}^{\prime\prime}\tan(\omega t)\,.$~~~ (2.17)
    We define the averaged transmission and reflection coefficients by performing the principle integration from $\frac{\pi}{2\omega}$ to $\frac{\pi}{2\omega}$,
[…]
@@ 2.1.1 The equations of perturbations
[…]
@@ 3.1 On the UV boundary condition for asymptotic AdS2 solution
where $A\not=0$ and $B\leq 0$. For a given profile of $Q$, the scalar field $\phi(z)$ and the potential $V(\phi)$ are determined from Eq. (2.5). The localized scalar operator in the dual ICFT has a conformal dimension of $1/2$ for $B<0$ and $1$ for $B=0$ while the coefficient of $\psi(z)$ in front of $z^{3}$ nonzero. Below, we show that under the sourceless boundary condition for the fluctuations, the relation $\mathcal{R}+\mathcal{T}=1$ holds, consistent with energy conservation as discussed in Sec. 2.1.
With the above asymptotic behavior, the solution to Eq. (2.22) has a UV expansion
$\xi(z)=\xi_{1}\,z+\xi_{2}\,z^{2}+\frac{1}{2}\left[-1-\mathcal{R}+\mathcal{T}-\omega^{2}(1+A^{2})\xi_{1}+\frac{2B\xi_{2}}{A}\right]z^{3}+\cdots\,$, (3.2)
where $\xi_{1}$ and $\xi_{2}$ are integration constants, formed by combining the homogeneous solution’s integration constants with the asymptotic expansion terms of the particular solution in (2.28).
@@
where $p_{1}$ and $p_{2}$ are two integral constants. Here, $p(z)$ in (3.3) represents the asymptotic expansion of the complete solution. Unlike our previous analysis of $\xi(z)$ above, the contribution from the sourced term of (2.23) to $p(z)$ only appears in higher-order terms of the expansion. For $B=0$ while the coefficient in front of $z^{3}$ nonzero, the UV expansion is
@@
where $p_{1}$ and $p_{2}$ are again integration constants. For $B<0$, applying the sourceless condition $p_{1}=0$ yields, $p(z)=p_{2}\sqrt{z}\left(1+\mathcal{O}(z)\right).$ For $B=0$ while the coefficient in front of $z^{3}$ nonzero, the sourceless condition gives $p(z)=p_{2}\,z+\cdots$. Note that in (3.1) if the nonzero higher order term of $z^{n}$ first starts with an integer $n$ of $n\geq 2$, the scalar fluctuation have higher conformal dimension (n1)/2(n-1)/2.
@@
when $z\to 0$ . By imposing the infalling boundary condition for $\xi_{l}$ near the IR and the above UV boundary condition, we can solve the system and obtain the final solution for $\xi_{l}$ and the transmission coefficient $\mathcal{T}$.
@@ 3.2 An example from numerical studies
where the interface entropy is defined in [21], and $c=3L/2G$ is the central charge of the left/right CFT. We restrict our discussion to the region $b>0$ to ensure a positive interface entropy. The UV expansion of this profile matches exactly with Eq. (3.1) upon noticing that $A=a$ and B=baB=b-a.
@@
In Fig. 2, we show the plots of $\mathcal{T}^{\prime}$ and $\phi_{t}$ as functions of $a$. In the upper two plots, we fix the frequency and tune $b$. Both $\mathcal{T}^{\prime}$ and $\phi_{t}$ exhibit oscillations with a period of approximately $\omega a/\langle\mathcal{O}\rangle^{2}\simeq 0.4$. The amplitude of these oscillations decreases as $a\omega/\langle\mathcal{O}\rangle^{2}$ increases for fixed $b$. Furthermore, the upper two plots in Fig. 2 show that the oscillation amplitude also decreases when $b$ is increased at a fixed value of $a\omega/\langle\mathcal{O}\rangle^{2}$.
Figure 3: The plots of $\mathcal{T}^{\prime}$ and $\phi_{t}$ as functions of $b$, for different values of $a$ with fixed $\omega/\langle\mathcal{O}\rangle^{2}=0.01$ (upper two plots) and for different values of $\omega/\langle\mathcal{O}\rangle^{2}$ with fixed $a=2$ (lower two plots). In the upper two plots, $a=1$ (black), 33 (blue), 55 (purple), and 77 (red).
In Fig. 3, we show the plots of $\mathcal{T}^{\prime}$ and $\phi_{t}$ as functions of $b$. Due to the null energy condition constraint $a\geq b$, we focus on the region where $b\in(0,1]$ with $a\geq 1$. We observe that, for the parameters considered, $\mathcal{T}^{\prime}$ and $\phi_{t}$ exhibit monotonic behavior for all values of $b$ within this region. This behavior contrasts with the trends observed in Fig. 2. Moreover, we find that when $b\to 0$, $\mathcal{T}^{\prime}$ approaches distinct values, which differs from its dependence on $a$. Furthermore, for a fixed value of $b$, we observe that $\mathcal{T}^{\prime}$ increases when $\omega/\langle\mathcal{O}\rangle^{2}$ decreases within the studied parameter regions, and this is consistent with the observations in Fig. 1 for the frequencies below the “dip-ramp” feature shown there. [^6] We have numerically verified that with a fixed $a=2$ and various $b$ values in the range $(0,1]$, the transmission coefficient $\mathcal{T}^{\prime}$ decreases monotonically as $\omega/\langle\mathcal{O}\rangle^{2}$ increases across the interval $[0.01,0.1]$.
3.3 An analytical example with UV and IR asymptotically AdS2
In this subsection we show a brane profile for which analytical solutions of the transmission coefficients can be derived. In this example, the induced metric on the interface brane is asymptotically AdS2 in both the UV and IR regions. From the analytical solution, we observe key features that were previously identified in the numerical example in Sec. 3.2. Another similar example is provided in the appendix C.
@@ 3.3 An analytical example with UV and IR asymptotically AdS2
The null energy condition requires that $ab>0$. Here we consider $a>0$ and $b>0$. The background brane geometry is asymptotically AdS with radius $L\sqrt{1+a^{2}}$ in the UV and radius $L$ in the IR. From the UV expansion, it exactly matches with (3.1) by identifying $A=a$ and B=ab/2B=-ab/2.
@@
[…]
@@
$U\left(\frac{1}{2},\ 1,\ 2i\sqrt{2}\frac{\omega}{b}\right)$ $\to-\frac{\gamma+\log\frac{i\omega}{\sqrt{2}b}}{\sqrt{\pi}}\,,~~~~\frac{\omega}{b}U\left(\frac{3}{2},\ 2,\ 2i\sqrt{2}\frac{\omega}{b}\right)\to-\frac{i}{\sqrt{2\pi}}\,$, (3.24)
@@ 3.4 Another kind of UV asymptotically AdS geometry
Unitarity of the dual system constraints $\xi_{3}\in[-1,0]$. For the specific case $\psi(z)=-az^{2}$, the exact solution for $\xi(z)$ exists. The infalling boundary condition and $\xi_{1}=0$ imply $\xi_{3}=0$, leading to $\mathcal{R}=0,\mathcal{T}=1$.
@@ A.2 Equations for the fluctuations
From Eqs. (A.1) (A.4) and (A.5), seven equations for the fluctuations can be obtained. Among them, only five equations are independent. More precisely, Eq. (A.5) and the $zz$ component of (A.4) can be derived from the other five equations. These five independent equations can be simplified as Eqs. (2.22), (2.23), (2.25), (2.26) and (2.27).
@@
[…]